what is the mass of chair inside a classroom where g=9.8 m/s2 if it weights 300 N? what it is weight if it were on the surface of the moon
1
Expert's answer
2014-02-25T08:57:24-0500
mass = Weight/g = 300/9.8 = 30.6 kg. On the Moon g(Moon) = 1.62 m/s^2, so Weight(Moon) = mass*g(Moon) = 49.6 N.